Bacterial cells grown for many generations in a "heavy" isotope of nitrogen are shifted to a "light" isotope of nitrogen. After three rounds of DNA replication, _______, _______, and ___________% of the DNA in the cells will have heavy, intermediate, and light density, respectively.

a. 0, 50, 50

b. 50, 50, 0

c. 25, 25, 50

d. 50, 0, 50

e. 0, 25, 75
Answer #1

the DNA with heavy nitrogen undergoes first round of replication to produce 2 DNA molecules with intermediate density, then these DNA undergoes second round of replication to produce 2 DNA molecules with intermediate density and 2 DNA molecules with light nitrogen, that is the number of DNA molecules with the intermediate density remains 2, after third round of replication the 2 DNA molecules with light nitrogen replicated to produce 4 DNA molecules with light nitrogen, and 2 DNA molecules with intermediate density, that is there are no bands with heavy nitrogen, and out of 8 double-stranded DNA molecules 2 have intermediate density and 6 are light. so

in the image the blue lines show DNA with light nitrogen, black lines show DNA strand with heavy nitrogen, if both strands are blue then the DNA has light density if one strand is heavy and another strand is light then the DNA has intermediate density, if both strands have heavy nitrogen the Double-stranded DNA is heavy.

percentage of light bands= (number of light bands/8)100

= (6/8)100

= 75%

so percentage of DNA molecules with intermediate density= 100-57%=25%

so the answer is e. 0, 25, 75